# Fuzz Driver Synthesis for Rust Generic APIs

### 17 Dec 2023 v1

@@ Abstract.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austriaconference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austriaccs: Software and its engineering Software defect analysisccs: Software and its engineering Automated static analysisccs: Security and privacy Software security engineering
@@ 1. Introduction
In this paper, we investigate library fuzzing for Rust crates, with a specific focus on addressing the automated fuzz driver synthesis problem for generic APIs. The problem is important because Rust is a system programming language that places a strong emphasis on security and robustness. Bugs within Rust libraries have the potential to impact all downstream code. Generic APIs are widely employed in Rust libraries, offering developers a viable means to avoid repetitively implementing the same API for different types. A generic API generally has one or multiple parameters of generic types e.g., T, which will be monomorphized to concrete types upon invocation. However, existing work on fuzz driver synthesis (e.g., RULF (Jiang et al., 2021)) often neglects generic APIs or treats them merely as non-generic APIs. As a result, it either falls short in generating valid fuzz drivers for generic APIs or fails to produce diversified monomorphic versions for the same generic API.
@@ 2.1. Generic Type and Generic API
Rust provides traits(official team, 2023e) to describe an abstract interface that types can implement. Trait can be regarded as a set of interface declarations. A type can implement a trait by implementing all interfaces declared in the trait. Based on the concept of trait, Rust provides a type constraint mechanism for the generic API called trait bounds(official team, 2023d). A type parameter can specify the trait bounds to require the concrete type for this type parameter must implement the specific traits. If the type for the type parameter unsatisfies the requirement, the compiler will report a compilation error. On the other hand, as the compiler only accepts the type implementing these traits, the developer can safely invoke the interface of the traits for this type parameter in generic API. Trait bounds in Rust are composable, meaning if you want to specify that a type parameter T must implement both trait A and B, you can use the syntax T:A + B to express this requirement. Rust also provides a ”where” clause, which allows programmers to declare trait bounds on any type or associated types(official team, 2023b). This feature empowers Rust programmers to write type constraints flexibly and expressively.
@@ 2.2.1. Generic API Dependency Graph
  • $P$ is the set of producer edges, where $C\subseteq N_{T}\times N_{A}$.
@@
Figure 3 demonstrates a generic API dependency graph for an example library. The library source code is shown in figure 2. The graph contains 2 non-generic nodes and 3 generic nodes. There are 2 concrete types Ty1, Ty2 and 2 generic types T and Vec<T>. The consumer and producer edges are constructed by the input types and return types for each API. Finally, there are two match edge(Ty1normal-$\rightarrow$ T, Ty2normal-$\rightarrow$ T) that represent both Ty1 and Ty2 can match with generic type T.
@@ 2.2.2. The Challenges of Fuzz Driver Synthesis
To synthesize valid fuzz drivers, we have to generate corresponding valid API sequences. A valid API sequence ensures that all APIs can obtain the instance of its input types from previous APIs or fuzzer so that the fuzz driver can be compiled successfully. However, the generic APIs in a valid API sequence should be replaced with specific monomorphic APIs so that we can determine the exact data dependencies. In other words, we should determine the concrete types for every type parameter in the generic API and ensure the type combination for type parameters can satisfy the trait bounds of the generic API.
@@ 5.1.2. Library Selection
We generate all fuzz drivers by our approach. As we mention in 4.3, all generated fuzz drivers contain at least one generic API. We build all fuzz drivers with command cargo afl build.
@@ 5.2.2. Coverage of our approach
Table 2 presents the coverage for API and generic API in each library, and shows our approach can achieve 0.7+ API coverage for 14 libraries and 0.6+ generic API coverage for 9 libraries. However, it does not work well for some libraries, the coverage of our tool varies among different libraries. Across all libraries, the API coverage spans from 0.01 to 0.96, and the generic API coverage ranges from 0.01 to 0.75. The total API coverage is 0.66, while the total generic API coverage is 0.27. To understand the culprits of the low generic API coverage, We study the library with the lowest generic API coverage (axum-core). We have 72 uncovered generic APIs, and we find that 68 of them are caused by using the types or traits of third-party libraries in the generic API signature in this library. Figure 9 shows an example of this situation. The axum-core declares a trait FromRequest with an interface from_request. However, the first argument of this interface is a third-party type http::Request. Since we lack the API information from http, we can not instantiate a http::Request. As a result, all APIs related to this trait are uncovered. This issue is difficult to resolve since we implement our tool based on rustdoc, which is designed for documenting a single library. It takes a lot of engineering effort to adopt our tool to extract the API information from third-party libraries. The reason for the rest of the 4 uncovered APIs is that there are no types that implement the trait required by the trait bounds of the generic API. Since we generate monomorphic API from all existing types in the library, our approach can not cover this situation.
@@ 5.4.1. Overview
We also examine all bug reports to ascertain how many bugs are related to the generic API. A bug is considered generic API-related if it is either caused by a generic API or requires the invocation of at least one generic API. Out of 23 bugs, we found 18 bugs are related to generic APIs.

### 19 Dec 2023 v2

@@ Abstract.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austriacopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: arxiv pre-print versions; ; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software defect analysisccs: Software and its engineering Automated static analysisccs: Security and privacy Software security engineering
@@ 1. Introduction
In this paper, we investigate library fuzzing for Rust libraries, with a specific focus on addressing the automated fuzz driver synthesis problem for generic APIs. The problem is important because Rust is a system programming language that places a strong emphasis on security and robustness. Bugs within Rust libraries have the potential to impact all downstream code. Generic APIs are widely employed in Rust libraries, offering developers a viable means to avoid repetitively implementing the same API for different types. A generic API generally has one or multiple parameters of generic types e.g., T, which will be monomorphized to concrete types upon invocation. However, existing work on fuzz driver synthesis (e.g., RULF (Jiang et al., 2021)) often neglects generic APIs or treats them merely as non-generic APIs. As a result, it either falls short in generating valid fuzz drivers for generic APIs or fails to produce diversified monomorphic versions for the same generic API.
@@ 2.1. Generic Type and Generic API
Rust provides trait(official team, 2023e) to describe an abstract interface that types can implement. Trait can be regarded as a set of interface declarations. A type can implement a trait by implementing all interfaces declared in the trait. Based on the concept of trait, Rust provides a type constraint mechanism for the generic API called trait bounds(official team, 2023d). A type parameter can specify the trait bounds to require the concrete type for this type parameter must implement the specific traits. If the type for the type parameter unsatisfies the requirement, the compiler will report a compilation error. On the other hand, as the compiler only accepts the type implementing these traits, the developer can safely invoke the interface of the traits for this type parameter in generic API. Trait bounds in Rust are composable, meaning if you want to specify that a type parameter T must implement both trait A and B, you can use the syntax T:A + B to express this requirement. Rust also provides a ”where” clause, which allows programmers to declare trait bounds on any type or associated types(official team, 2023b). This feature empowers Rust programmers to write type constraints flexibly and expressively.
@@ 2.2.1. Generic API Dependency Graph
  • $P$ is the set of producer edges, where $P\subseteq N_{T}\times N_{A}$.
@@
Figure 3 demonstrates a generic API dependency graph for an example library. The library source code is shown in figure 2. The graph contains 2 non-generic nodes and 3 generic nodes. There are 2 concrete types Ty1, Ty2 and 2 generic types T and Vec<T>. The consumer and producer edges are constructed by the input types and return types for each API. Finally, there are two match edges(Ty1normal-$\rightarrow$ T, Ty2normal-$\rightarrow$ T) that represent both Ty1 and Ty2 can match with generic type T.
@@ 2.2.2. The Challenges of Fuzz Driver Synthesis
To synthesize valid fuzz drivers, we have to generate corresponding valid API sequences. A valid API sequence ensures that all APIs can obtain instances of their input types from previous APIs or the fuzzer so that the fuzz driver can be compiled successfully. However, the generic APIs in a valid API sequence should be replaced with specific monomorphic APIs so that we can determine the exact data dependencies. In other words, we should determine the concrete types for every type parameter in the generic API and ensure the type combination for type parameters can satisfy the trait bounds of the generic API.
@@ 5.1.2. Library Selection
We generate all fuzz drivers by our approach. As we mention in section  4.3, all generated fuzz drivers contain at least one generic API. We build all fuzz drivers with the command cargo afl build.
@@ 5.2.2. Coverage of our approach
Table 2 presents the coverage for API and generic API in each library. Our approach can achieve 0.7+ API coverage for 14 libraries and 0.6+ generic API coverage for libraries. However, it does not work well for some libraries, and the coverage of our tool varies among different libraries. Across all libraries, the API coverage spans from 0.01 to 0.96, and the generic API coverage ranges from 0.01 to 0.75. The total API coverage is 0.66, while the total generic API coverage is 0.27. To understand the culprits of the low generic API coverage, We study the library with the lowest generic API coverage (axum-core). We have 72 uncovered generic APIs, and we find that 68 of them are caused by using the types or traits of third-party libraries in the generic API signature in this library. Figure 9 shows an example of this situation. The axum-core declares a trait FromRequest with an interface from_request. However, the first argument of this interface is a third-party type http::Request. Since we lack the API information from http, we can not instantiate a http::Request. As a result, all APIs related to this trait are uncovered. This issue is difficult to resolve since we implement our tool based on rustdoc, which is designed for documenting a single library. It takes a lot of engineering effort to adopt our tool to extract the API information from third-party libraries. The reason for the rest of the 4 uncovered APIs is that there are no types that implement the trait required by the trait bounds of the generic API. Since we generate monomorphic API from all existing types in the library, our approach can not cover this situation.
@@ 5.4.1. Overview
We also examine all bugs to ascertain how many bugs are related to the generic API. A bug is considered generic API-related if it is either caused by a generic API or requires the invocation of at least one generic API. Out of 23 bugs, we found 18 bugs are related to generic APIs.